# REvolve: Reward Evolution with Large Language Models for Autonomous Driving

### v1

@@ B.1 REvolve Fitness
As state in § 3.3, REvolve employs a user interface to gather human feedback on behavior $t$ sampled from a trained policy. Human evaluators review pairs of videos, submitting their preferences. A fitness function then maps this data to a real-valued fitness score and a natural language feedback i.e. $(\sigma,\lambda)\coloneqq F(t)$. We start by introducing the Elo rating system that maps the preference data to a fitness score. For two reward function individuals $A$ and $B$ with fitness scores $\sigma_{A}$ and $\sigma_{B}$ respectively, the expected scores $E_{A}$ and $E_{B}$ are calculated using the logistic function:
@@
The process greatly simplifies the feedback interface, as participants only need to choose their preferred video from a pair, rather than providing an absolute rating. Additionally, this method reduces individual biases that could emerge from various factors, such as the need to adjust ratings at different stages of evolution – from the initial phase, where many policies underperform, to later stages, when performances improve. To overcome this, we compare individuals not only within the same generation but also across different generations, updating the fitness scores of all individuals at the end of each generation [^4] This approach is akin to comparing chess grandmasters Gary Kasparov and Magnus Carlsen, each a dominant player in his respective era. However, a direct comparison would require them to be matched in their primes; otherwise, the basis of comparison would be skewed and unjustified.. In direct rating-based systems, participants must repeatedly rate all generations, adjusting their scales each time. In contrast, our preference-based rating system allows us to efficiently utilize all existing preference data across generations to assign new fitness scores, thus requiring minimal data collection for the current generation.
B.2 Eureka Automatic Feedback
Manually Designed Fitness Function.
@@ B.2 Eureka Automatic Feedback
$\text{TotalFitness}(collision,v,d)=\begin{cases}\text{collision penalty},&\text{if collision occurs}\\ \text{speed penalty},&\text{if }v<v_{\text{min\_limit}}\text{ or }v>v_{\text{max\_limit}}\\ \text{distance penalty},&\text{if }d>d_{\text{fail}}\\ \text{score}(v,d),&\text{otherwise}\end{cases}$ (2)
@@
with $d_{max}=0.5$m being the maximum tolerance for deviation from the waypoint. Both functions are shown in 8.
Figure 8: Visual representation of speed and distance scores.
In table 3, we summarize the fitness function components used for evaluating RL models.
Table 3: Fitness Function Hyperparameters Details
@@
Following Eureka, we track the statistics of all reward components at intermediate policy checkpoints throughout training. We refer the interested reader to the original paper [30].
@@ Appendix C Expert Designed Rewards
Following the methodologies advocated in [44, 52], the total reward at each time step, $R_{\text{total}}$, is constructed from a composite of critical driving behaviors. This section describes three configurations of speed rewards, each designed to optimize the agent’s driving performance by emphasizing speed consistency, smooth navigation, and collision avoidance.
Human Engineered:
Total Reward:
$r_{\text{total}}^{HE}=c_{\text{pos}}\times r_{\text{pos}}+c_{\text{sm}}\times r_{\text{sm}}+c_{\text{speed}}\times r_{\text{speed}}+c_{\text{sensor}}\times r_{\text{sensor}}$
Position Reward:
$r_{\text{pos}}=e^{-\delta\times((\text{min\_pos})^{2}-\beta)}$
Speed Reward:
$r_{\text{speed}}(\text{speed})=\begin{cases}-\frac{|\text{speed}-v_{\text{min}}|}{v_{\text{min}}}&\text{if speed}<v_{\text{min}}\\ -\frac{|\text{speed}-v_{\text{max}}|}{\text{speed}}&\text{if speed}>v_{\text{max}}\\ 1&\text{otherwise}\end{cases}$
Sensor Reward:
$r_{\text{sensor}}=\begin{cases}-\frac{(dis_{safe}-\text{distance})}{dis_{safe}}&\text{if distance}<dis_{safe}\\ 0.5&\text{otherwise}\end{cases}$
Smoothness Reward:
[…]
Table 4: Values and Corresponding Fitness Scores
@@
Table 5: Reward Components Hyperparameters Details
@@
Human Engineered Tuned:
@@
Total Reward:
$r_{\text{total}}^{HET}=c_{\text{pos}}^{opt}\times r_{\text{pos}}+c_{\text{sm}}^{opt}\times r_{\text{sm}}+c_{\text{speed}}^{opt}\times r_{\text{speed}}+c_{\text{sensor}}^{opt}\times r_{\text{sensor}}$
@@
Appendix D Prompts
@@
System Prompt In Box D that outlines the task description and expected input from the environment, complete with formatting tips to help guide GPT-4 in generating suitable reward functions.
@@
Abstracted Environment Variables In Box D that provides a list of (abstracted) observation and action variables in the environment, formatted as a Python class. This informs GPT-4 of variables it can use in reward design.
@@
Operator Prompts like crossover (Box D) and mutation (Box D). Additionally, we provide prompts for REvolve Auto baseline (Boxes D D).
Box 1: System Prompt You are a reward engineer trying to write a reward function to create an autonomous car agent to learn to navigate the environment. Your goal is to write a reward function for the environment that will help the agent learn the task described in text.
Task description:
The objective is to create an autonomous car agent to learn to navigate the environment. Specifically, the performance metric for the learning agent is to successfully drive without crashing, while maintaining smooth driving behavior and a speed between 9.0-10.5 m/s. The episode terminates successfully if the agent drives for 1000 consecutive time steps without crashing. Otherwise, the episode terminates in failure. The controlled actions are the steering and the speed. The car must avoid staying inactive and stay as close to the center of the road as possible to ensure optimal performance.
@@
Your reward function should use useful variables from the environment as inputs. An example reward function signature can be: ```python def compute_reward(object_pos: type, goal_pos: type) -> Tuple[type, Dict[str, type]]: ... return reward, {} ``` The output of the reward function should consist of two items: 1. The total reward, 2. A dictionary of each individual reward component. The code output should be formatted as a Python code string: "“‘python … “‘". 1. You may find it helpful to normalize to a fixed reward range like -1,1 or 0,1 by applying transformations like np.exp() to the overall reward and/or its reward components. 2. If you choose to transform a reward component, then you must introduce a temperature parameter inside the transformation function. This parameter must be a named variable in the reward function and it must not be an input variable. 3. Make sure the type of each input variable is correctly specified. 4. Most importantly, the reward’s code input variables must contain only attributes of the environment. Under no circumstance can you introduce a new input variable and assume variables that are not available. 5. Your output must always be the def function code with the output being the total reward. 6. Make sure you don’t give contradictory reward components.
@@
Box 2: Abstracted Environment Variables (as a Python class) from scipy.spatial.transform import Rotation as R import numpy as np class Env: def __init__(self): self.client = CarClient() action = [CarControls().steering, CarControls().throttle] def state(self): position_info = self.client.getCarState().kinematics_estimated curr_x = position_info.position.x_val curr_y = position_info.position.y_val yaw, pitch = self.get_orientation(position_info) linear_velocity = position_info.linear_velocity vx, vy = linear_velocity.x_val, linear_velocity.y_val angular_velocity_x, angular_velocity_y, angular_velocity_z = position_info.kinematics_estimated.angular_velocity.x_val, position_info.kinematics_estimated.angular_velocity.y_val, position_info.kinematics_estimated.angular_velocity.z_val speed = position_info.speed collision = self.client.simGetCollisionInfo().has_collided action_list = [steering_value1, steering_value2, steering_value3, steering_value4] min_pos = get_min_pos(self, position_info) total_step_counter = 0 episode_step_counter = 0 distance = 20 # default max range fo sensor is 20 def get_orientation(self, position_info): quaternion = np.array([position_info.orientation.w_val, position_info.orientation.x_val, position_info.orientation.y_val, position_info.orientation.z_val]) euler_angles = R.from_quat(quaternion).as_euler(’zyx’, degrees=True) return euler_angles[0], euler_angles[1] def get_min_pos(self, position_info): min_dis = float(’inf’) for i in data: stx, sty = i[0], i[1] dis = eucl_dis(stx, sty, curr_x, curr_y) if dis < min_dis: min_dis = dis return min_dis def calculate_front_distance(self): distance_data = self.client.getDistanceSensorData(vehicle_name="Car") distance = distance_data.distance # max sensor distance is 20
@@
Box 3: REvolve Crossover Prompt You are given a set of reward functions used to train autonomous driving agents. The reward functions have fitness scores that reflect the agent’s driving performance – higher scores indicate superior driving behavior. For each reward function, we collected human feedback on what aspects of the agent’s driving are satisfactory and what aspects need improvement. Additionally, we tracked the values of the individual components in the reward function after every <episodes> episodes and noted the maximum, mean, minimum values observed.  <EXAMPLES> Your task: Combine high-performing reward components from different reward functions based on human feedback and the data tracked, to enhance the agent’s driving capability further. Clearly explain how you intend to combine them and your rationale behind improving the driving performance. Write the combined reward function code. Output of the reward function should consist of: 1: The total reward. 2: A dictionary of each individual reward component. Code output should be formatted as a Python code string: ```python``` Tips for reward function design: 1. Normalize to a fixed reward range like -1,1 or 0,1 by applying transformations like np.exp() to the overall reward or its components. 2. If transforming a reward component, introduce a temperature parameter inside the transformation function. This parameter must be a named variable in the reward function and it must not be an input variable. Each transformed reward component should have its own temperature variable and you must carefully set its value based on its effect in the overall reward. 3. Ensure the type of each input variable is correctly specified. 4. The reward’s code input variables must contain only attributes of the environment. Under no circumstance can you introduce a new input variable. 5. Always output the def function code with the output being the total reward. 6. Avoid contradictory reward components. 7. Ensure that the penalties for collision and inactivity are given precedence over other reward components. This means that in the event of a collision or the car staying still, the total reward for that time step should primarily reflect the penalty for these specific conditions, potentially negating other positive rewards accrued from different aspects of the driving performance.
Box 4: REvolve Mutation Prompt You are given a reward function used to train autonomous driving agents. The reward functions has a fitness scores that reflects the agent’s driving performance – higher scores indicate superior driving behavior. We also collected human feedback on what aspects of the agent’s driving are satisfactory and what aspects need improvement. Additionally, we tracked the values of the individual components in the reward function after every <episodes> episodes and noted the maximum, mean, minimum values observed.
<EXAMPLES>
Your task is to iterate on the reward function by mutating a single component to enhance the agent’s driving capability further. Some helpful tips for analyzing the reward components: If the values for a certain reward component are nearly identical throughout, then this means the training is not able to optimize this component as it is written. You may consider:
@@
1. Rescaling it to a proper range or the value of its temperature parameter 2. Re-writing the reward component 3. Discarding the reward component The mutation process involves the following steps: 1. First, select a reward component based on the given guidelines, human feedback, and the values of individual reward components. 2. Next, clearly explain how you intend to mutate the selected component and your rationale behind improving the driving performance. This could involve adjusting its scale, rewriting its formula, or other modifications. 3. Finally, write the mutated reward function code.
@@
Box 5: REvolve Auto Crossover Prompt You are given a set of reward functions used to train autonomous driving agents. The reward functions have fitness scores that reflect the agent’s driving performance – higher scores indicate superior driving behavior. We also tracked the values of the individual components in the reward function after every <episodes> episodes and noted the maximum, mean, minimum values observed.
<EXAMPLES>
Your task: Combine high-performing reward components from different reward functions to enhance the agent’s driving capability further. Clearly explain how you intend to combine them and your rationale behind improving the driving performance. Write the combined reward function code.
@@
Box 6: REvolve Auto Mutation Prompt You are given a reward function used to train autonomous driving agents. The reward function has fitness scores that reflect the agent’s driving performance – higher scores indicate superior driving behavior. We also tracked the values of the individual components in the reward function after every <episodes> episodes and noted the maximum, mean, and minimum values observed.
<EXAMPLES>
Your task is to iterate on the reward function by mutating a single component to enhance the agent’s driving capability further. Some helpful tips for analyzing the reward components: If the values for a certain reward component are nearly identical throughout, then this means the training is not able to optimize this component as it is written. You may consider:
1. Rescaling it to a proper range or the value of its temperature parameter 2. Re-writing the reward component 3. Discarding the reward component
The mutation process involves the following steps:
@@
1. First, select a reward component based on the given guidelines. 2. Next, clearly explain how you intend to mutate the selected component and your rationale behind improving the driving performance. This could involve adjusting its scale, rewriting its formula, or other modifications. 3. Finally, write the mutated reward function code.
@@
The output of the reward function should consist of two items:
1. The total reward, 2. A dictionary of each individual reward component.
The code output should be formatted as a Python code string: ```python``` Some helpful tips for writing the reward function code:

### v2

@@ B.1 REvolve Fitness
As stated in Section 3.3, REvolve employs a user interface to gather human feedback on behavior $t$ sampled from a trained policy. Human evaluators review pairs of videos, submitting their preferences. A fitness function then maps this data to a real-valued fitness score and a natural language feedback i.e. $(\sigma,\lambda)\coloneqq F(t)$. We start by introducing the Elo rating system that maps the preference data to a fitness score. For two reward function individuals $A$ and $B$ with fitness scores $\sigma_{A}$ and $\sigma_{B}$ respectively, the expected scores $E_{A}$ and $E_{B}$ are calculated using the logistic function:
@@
The process greatly simplifies the feedback interface, as participants only need to choose their preferred video from a pair rather than providing an absolute rating. Additionally, this method reduces individual biases that could emerge from various factors, such as the need to adjust ratings at different stages of evolution – from the initial phase, where many policies underperform, to later stages, when performances improve. To overcome this, we compare individuals not only within the same generation but also across different generations, updating the fitness scores of all individuals at the end of each generation [^5] This approach is akin to comparing chess grandmasters Gary Kasparov and Magnus Carlsen, each a dominant player in his respective era. However, a direct comparison would require them to be matched in their primes; otherwise, the basis of comparison would be skewed and unjustified.. In direct rating-based systems, participants must repeatedly rate all generations, adjusting their scales each time. In contrast, our preference-based rating system allows us to efficiently utilize all existing preference data across generations to assign new fitness scores, thus requiring minimal data collection for the current generation.
B.2 Automatic Feedback
Autonomous Driving.
@@ B.2 Automatic Feedback
$\sigma=\begin{cases}\text{collision penalty},&\text{if collision occurs}\\ \text{speed penalty},&\text{if }v<v_{\text{min\_limit}}\text{ or }v>v_{\text{max\_limit}}\\ \text{distance penalty},&\text{if }d>d_{\text{fail}}\\ \text{score}(v,d),&\text{otherwise,}\end{cases}$ (2)
@@
with $d_{max}=0.5$m being the maximum tolerance for deviation from the waypoint. Both functions are shown in Figure 9. In Table 5, we summarize the fitness function components used for evaluating RL models.
Table 5: Fitness Function Hyperparameters used for calculating the fitness score for AirSim environment for Autonomous Driving task.
@@
Figure 9: Visual representation of speed and distance fitness scores for autonomous driving in the AirSim environment.
Humanoid Locomotion.
The fitness function for the Humanoid environment evaluates the agent based on its ability to maintain forward locomotion without falling over an episode of $T_{\text{max}}=1000$ steps.
$\sigma=\begin{cases}\frac{1}{T_{\text{max}}}\sum_{t=1}^{T_{\text{max}}}v_{t}^{x},&\text{if }T=T_{\text{max}}\\ 0,&\text{if }T<T_{\text{max}}\end{cases}$ (3)
where, $v_{t}^{x}$ is the forward velocity along the x-axis at time step $t$. $T$ is the total number of steps the agent remained upright. This function rewards agents that achieve higher forward velocities while also surviving the entire episode.
Adroit Hand.
The fitness function for the Adroit Hand environment rewards agents based on the number of steps taken to successful task completion. As shown in Figure 10, completing the task in fewer steps results in a higher fitness score $\in[0.5,1.0]$. If the task is not completed, the fitness score is 00.
Let $steps$ be the number of steps taken to complete the task ($50\leq steps\leq 400$), $T_{\text{min}}=50$ be the minimum possible steps, $T_{\text{max}}=400$ be the maximum allowed steps, $a=-\dfrac{1}{700}$, $b=\dfrac{75}{70}$ are constants defining the linear relationship. The fitness score $\sigma$ is defined as:
$\sigma=\begin{cases}a\times steps+b,&\text{if the task is successfully completed}\\ 0,&\text{if the task is not completed}\end{cases}$ (4)
This score measures the efficiency of the Adroit Hand in completing the task, awarding a perfect score of 1 for completing it in 50 steps or fewer, with a linear decrease to 0.5 for completing the task in the maximum allowable 400 steps, and a score of 0 if the task is not completed.
Figure 10: Visual representation of fitness score for Adroit Hand when completing the task.
@@
Following Eureka, we track the statistics of all reward components at intermediate policy checkpoints throughout training. We refer the interested reader to the original paper (Ma et al., 2024a).
@@ Appendix C Expert Designed Rewards
C.1 Autonomous Driving
Following the methodologies advocated in Spryn et al. (2018), the total reward at each time step, $R_{\text{total}}$, is constructed from a weighted sum of speed consistency, smooth navigation, and collision avoidance behavior of the agent. In what follows, we describe how the weights were calculated.
Table 6: Grid Search Weights: Values and Corresponding Fitness Scores. The quadruple of weights yielding the highest fitness score is highlighted in bold.
@@ C.1 Autonomous Driving
Table 7: Hyperparameters used in all reward components designed by Expert in AirSim in Autonomous Driving task.
@@
Human Expert Designed (HED):
$r_{\text{total}}^{\text{HED}}=c_{\text{pos}}\times r_{\text{pos}}+c_{\text{sm}}\times r_{\text{sm}}+c_{\text{speed}}\times r_{\text{speed}}+c_{\text{sensor}}\times r_{\text{sensor}}$
Position Reward:
$r_{\text{pos}}=e^{-\delta\times((\text{min\_pos})^{2}-\beta)}$
Speed Reward:
$r_{\text{speed}}=\begin{cases}-\frac{|\text{speed}-v_{\text{min}}|}{v_{\text{min}}}&\text{if speed}<v_{\text{min}}\\ -\frac{|\text{speed}-v_{\text{max}}|}{\text{speed}}&\text{if speed}>v_{\text{max}}\\ 1&\text{otherwise}\end{cases}$
Sensor Reward:
$r_{\text{sensor}}=\begin{cases}-\frac{(dis_{\text{safe}}-\text{distance})}{dis_{\text{safe}}}&\text{if distance}<dis_{\text{safe}}\\ 0.5&\text{otherwise}\end{cases}$
Smoothness Reward:
[…]
Human Expert Designed Tuned (HEDT):
@@
$r_{\text{total}}^{\text{HEDT}}=c_{\text{pos}}^{opt}\times r_{\text{pos}}+c_{\text{sm}}^{opt}\times r_{\text{sm}}+c_{\text{speed}}^{opt}\times r_{\text{speed}}+c_{\text{sensor}}^{opt}\times r_{\text{sensor}}$
@@
C.2 Humanoid Locomotion
In the humanoid environment, the agent is tasked with moving forward while maintaining a healthy state. The reward function is designed to encourage forward movement, staying upright, and minimizing control effort. The total reward $R$ is computed as:
$R=R_{\text{forward}}+R_{\text{healthy}}-R_{\text{control}}$
Where:
  • $R_{\text{forward}}=2.5\times w_{fwd}\times x\_velocity$
    This term provides a reward proportional to the agent’s forward velocity, incentivizing fast forward movement, with $w_{fwd}=1.25$.
  • $R_{\text{healthy}}$ is a reward for maintaining a healthy state, where the agent’s height is within a predefined range:
    $R_{\text{healthy}}=\begin{cases}15&\text{if the agent is healthy}\\ 0&\text{if the agent is unhealthy}\end{cases}$
  • $R_{\text{control}}=0.5\times w_{\text{ctrl}}\times\|\mathbf{\text{action}}\|_{2}^{2}$
    A penalty is applied based on the magnitude of the agent’s actions, discouraging excessive or inefficient control efforts with $w_{\text{ctrl}}=0.1$.
The total reward encourages the agent to move forward efficiently while staying upright and minimizing unnecessary control actions. The episode terminates if the agent becomes unhealthy (e.g., by falling) or after a maximum number of steps.
C.3 Adroit Hand
In the Adroit Hand Environment, the agent’s task is to manipulate a door handle and open a door. The reward function is designed to incentivize both reaching the handle and successfully opening the door. The total reward $R$ is a combination of multiple components:
$R=R_{\text{handle}}+R_{\text{door}}+R_{\text{velocity}}+R_{\text{bonus}}$
Where:
[…]
    This term rewards the agent for minimizing the distance between its palm and the door handle.
  • $R_{\text{door}}=-0.1\times(\text{goal\_distance}-1.57)^{2}$
    This component provides a reward for moving the door hinge closer to the desired goal position (typically when the door is open).
  • $R_{\text{velocity}}=-1\times 10^{-5}\times\sum\text{qvel}^{2}$
    A penalty is applied for excessive joint velocities (qvel), encouraging smooth and efficient movements.
  • $R_{\text{bonus}}$ is a set of bonus rewards based on how much the door has been opened:
    $R_{\text{bonus}}=\begin{cases}+2&\text{if }\text{goal\_distance}>0.2\\ +8&\text{if }\text{goal\_distance}>1.0\\ +10&\text{if }\text{goal\_distance}>1.35\ (\text{goal achieved})\end{cases}$
[…]
Appendix D REvolve Cost Analysis & Comparisons
Figure 11: Comparison of REvolve with existing frameworks used to train autonomous driving agents. (a.) Expert Designed reward function $R$ to train the autonomous driving agent; (b.) learning from demonstrations (LfD) and learning from interventions (LfI), where the agent is trained to imitate the human; (c.) humans acting as reward functions within the training loop, by assessing policy rollouts and providing scalar rewards, which requires significant manual effort; (d.) Reinforcement learning from human feedback (RLHF), where human preference data is used to train an additional (black-box) reward model. This reward model is used as a proxy for human rewards to train the autonomous driving policy; (e) Proposed REvolve, which uses GPT-4 as a reward function generator $G$ and evolves them based on (minimal) human feedback on the rollouts sampled from the trained models. This feedback is directly incorporated into the reward design process. REvolve outputs interpretable reward functions, thereby avoiding learning an additional reward model. Here, $\pi\in\Pi$ is a trainable policy in the set of policies $\Pi$. The trainable blocks are denoted by the symbol of the flame.
Computation cost: REvolve = Eureka. Both REvolve and Eureka output 16 reward functions and train 16 policies in each generation.
Data requirement: RLHF >>much-greater-than$>$ > REvolve: RLHF requires massive amounts of curated data to train a reward model. Note that the input to a reward model is a video (rollout of a policy), and the output is a preference score, requiring a heavy neural model. As a reference, the common video classification model I3D trains on the Kinetics dataset, which has 650,000 video clips (Carreira & Zisserman, 2017). IRL >>much-greater-than$>$ > REvolve: Similarly, Inverse Reinforcement Learning requires human demonstrations and interventions. In contrast, REvolve uses 10 human evaluators per generation, each assessing 20 data samples (video pairs) – hence, for 5 generations $10\times 20\times 5=1000$ curated samples only. Compare this to the Lyft self-driving dataset (Houston et al., 2021) used for motion prediction which has more than 1000 hours of data collected over 4 months by 20 cars.
Storage Cost: REvolve $>$ Eureka. The slightly higher cost is due to maintaining a population of best candidates rather than one candidate like Eureka. However, compared to the LLM memory cost (e.g., the biggest open-source LLM Llama 3.1 takes up 810 GB of memory), the storage cost is negligible – for every generation, the storage increases by  2.5GB (which includes stored policy weights and other metadata) – 12.5 GB after 5 generations. Also, REvolve does not require expert feedback.
Comparison with existing approaches in Reinforcement Learning and Human Feedback Due to its complex and hard-to-specify goals (Knox et al., 2023), autonomous driving presents a challenging problem for reinforcement learning. Autonomous vehicles must maneuver through complex environments and interact with other road users, all the while strictly adhering to traffic regulations. Common techniques to overcome these challenges include learning from demonstrations (LfD) (Schaal, 1996) or learning from intervention (LfI) (Saunders et al., 2018). However, they do not always ensure that the learned behaviors align with human values such as safety, robustness, and efficiency (Hadfield-Menell et al., 2016). An alternative is integrating human feedback by using humans as the reward function in the learning process directly (Thomaz & Breazeal, 2008; Knox & Stone, 2009). However, this is rather resource-intensive as each rollout needs to be evaluated by a human. REvolve directly integrates human feedback (on the trained policy) into the reward design process. Instead of using humans as reward functions directly, human values, preferences, and knowledge can also be incorporated via so-called human-in-the-loop (HITL) RL. Several techniques exist, for instance, human assessment (Thomaz & Breazeal, 2008; Christiano et al., 2017; Knox & Stone, 2009), human demonstration (Schaal, 1996), and human intervention (Saunders et al., 2018; Wang et al., 2018). They are particularly beneficial in scenarios like autonomous driving, where the reward function is challenging, and the nature of the desired outcome is complex. More recently, reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Rafailov et al., 2023) has been used successfully in training models like ChatGPT (OpenAI, 2022), providing guidance and alignment from user preferences. A further comparison of autonomous driving approaches with and human feedback is presented in Figure 11, which illustrates how REvolve uses human feedback to align learned policies with human driving standards without requirements for extra reward models, such as in the case of RLHF.
Appendix E Prompts
@@
System Prompt Boxes EEE outline the system prompts (i.e. task description and expected input from the environment, complete with formatting tips to help guide GPT-4) for generating suitable reward functions for Autonomous Driving, Humanoid Locomotion, and Adroit Hand, respectively.
@@
Abstracted Environment Variables Boxes EEE provides a list of (abstracted) observation and action variables in the environment, formatted as a Python class, for all three environments. This informs GPT-4 of variables it can use in reward design.
@@
Operator Prompts like crossover (Box E) and mutation (Box E). Additionally, we provide prompts for REvolve Auto baseline (Boxes EE).
@@
Box 1: System Prompt: Autonomous Driving You are a reward engineer trying to write a reward function to create an autonomous car agent to learn to navigate the environment. Your goal is to write a reward function for the environment that will help the agent learn the task described in text. Task description: The objective is to create an autonomous car agent to learn to navigate the environment. Specifically, the performance metric for the learning agent is to successfully drive without crashing, while maintaining smooth driving behavior and a speed between 9.0-10.5 m/s. The episode terminates successfully if the agent drives for 1000 consecutive time steps without crashing. Otherwise, the episode terminates in failure. The controlled actions are the steering and the speed. The car must avoid staying inactive and stay as close to the center of the road as possible to ensure optimal performance. Your reward function should use useful variables from the environment as inputs. An example reward function signature can be: ```python def compute_reward(object_pos: type, goal_pos: type) -> Tuple[type, Dict[str, type]]: ... return reward, {} ``` The output of the reward function should consist of two items: 1. The total reward, 2. A dictionary of each individual reward component. The code output should be formatted as a Python code string: ”“‘python … “‘”. 1. You may find it helpful to normalize to a fixed reward range like -1,1 or 0,1 by applying transformations like np.exp() to the overall reward and/or its reward components. 2. If you choose to transform a reward component, then you must introduce a temperature parameter inside the transformation function. This parameter must be a named variable in the reward function and it must not be an input variable. 3. Make sure the type of each input variable is correctly specified. 4. Most importantly, the reward’s code input variables must contain only attributes of the environment. Under no circumstance can you introduce a new input variable and assume variables that are not available. 5. Your output must always be the def function code with the output being the total reward. 6. Make sure you don’t give contradictory reward components.
@@
Box 2: Abstracted Environment Variables (as a Python class): Autonomous Driving from scipy.spatial.transform import Rotation as R import numpy as np class Env: def __init__(self): self.client = CarClient() action = [CarControls().steering, CarControls().throttle] def state(self): position_info = self.client.getCarState().kinematics_estimated curr_x = position_info.position.x_val curr_y = position_info.position.y_val yaw, pitch = self.get_orientation(position_info) linear_velocity = position_info.linear_velocity vx, vy = linear_velocity.x_val, linear_velocity.y_val angular_velocity_x, angular_velocity_y, angular_velocity_z = position_info.kinematics_estimated.angular_velocity.x_val, position_info.kinematics_estimated.angular_velocity.y_val, position_info.kinematics_estimated.angular_velocity.z_val speed = position_info.speed collision = self.client.simGetCollisionInfo().has_collided action_list = [steering_value1, steering_value2, steering_value3, steering_value4] min_pos = get_min_pos(self, position_info) total_step_counter = 0 episode_step_counter = 0 distance = 20 # default max range fo sensor is 20 def get_orientation(self, position_info): quaternion = np.array([position_info.orientation.w_val, position_info.orientation.x_val, position_info.orientation.y_val, position_info.orientation.z_val]) euler_angles = R.from_quat(quaternion).as_euler(’zyx’, degrees=True) return euler_angles[0], euler_angles[1] def get_min_pos(self, position_info): min_dis = float(’inf’) for i in data: stx, sty = i[0], i[1] dis = eucl_dis(stx, sty, curr_x, curr_y) if dis < min_dis: min_dis = dis return min_dis def calculate_front_distance(self): distance_data = self.client.getDistanceSensorData(vehicle_name="Car") distance = distance_data.distance # max sensor distance is 20
@@
Box 3: REvolve Crossover Prompt: Autonomous Driving You are given a set of reward functions used to train autonomous driving agents. The reward functions have fitness scores that reflect the agent’s driving performance – higher scores indicate superior driving behavior. For each reward function, we collected human feedback on what aspects of the agent’s driving are satisfactory and what aspects need improvement. Additionally, we tracked the values of the individual components in the reward function after every <episodes> episodes and noted the maximum, mean, minimum values observed.  <EXAMPLES> Your task: Combine high-performing reward components from different reward functions based on human feedback and the data tracked, to enhance the agent’s driving capability further. Clearly explain how you intend to combine them and your rationale behind improving the driving performance. Write the combined reward function code. Output of the reward function should consist of: 1: The total reward. 2: A dictionary of each individual reward component. Code output should be formatted as a Python code string: ```python``` Tips for reward function design: 1. Normalize to a fixed reward range like -1,1 or 0,1 by applying transformations like np.exp() to the overall reward or its components. 2. If transforming a reward component, introduce a temperature parameter inside the transformation function. This parameter must be a named variable in the reward function and it must not be an input variable. Each transformed reward component should have its own temperature variable and you must carefully set its value based on its effect in the overall reward. 3. Ensure the type of each input variable is correctly specified. 4. The reward’s code input variables must contain only attributes of the environment. Under no circumstance can you introduce a new input variable. 5. Always output the def function code with the output being the total reward. 6. Avoid contradictory reward components.
@@
Box 4: REvolve Mutation Prompt: Autonomous Driving You are given a reward function used to train autonomous driving agents. The reward functions has a fitness scores that reflects the agent’s driving performance – higher scores indicate superior driving behavior. We also collected human feedback on what aspects of the agent’s driving are satisfactory and what aspects need improvement. Additionally, we tracked the values of the individual components in the reward function after every <episodes> episodes and noted the maximum, mean, minimum values observed. <EXAMPLES> Your task is to iterate on the reward function by mutating a single component to enhance the agent’s driving capability further. Some helpful tips for analyzing the reward components: If the values for a certain reward component are nearly identical throughout, then this means the training is not able to optimize this component as it is written. You may consider: 1. Rescaling it to a proper range or the value of its temperature parameter 2. Re-writing the reward component 3. Discarding the reward component The mutation process involves the following steps: 1. First, select a reward component based on the given guidelines, human feedback, and the values of individual reward components. 2. Next, clearly explain how you intend to mutate the selected component and your rationale behind improving the driving performance. This could involve adjusting its scale, rewriting its formula, or other modifications. 3. Finally, write the mutated reward function code.
@@
Box 5: REvolve Auto Crossover Prompt: Autonomous Driving You are given a set of reward functions used to train autonomous driving agents. The reward functions have fitness scores that reflect the agent’s driving performance – higher scores indicate superior driving behavior. We also tracked the values of the individual components in the reward function after every <episodes> episodes and noted the maximum, mean, minimum values observed. <EXAMPLES> Your task: Combine high-performing reward components from different reward functions to enhance the agent’s driving capability further. Clearly explain how you intend to combine them and your rationale behind improving the driving performance. Write the combined reward function code.
@@
Box 6: REvolve Auto Mutation Prompt: Autonomous Driving You are given a reward function used to train autonomous driving agents. The reward function has fitness scores that reflect the agent’s driving performance – higher scores indicate superior driving behavior. We also tracked the values of the individual components in the reward function after every <episodes> episodes and noted the maximum, mean, and minimum values observed. <EXAMPLES> Your task is to iterate on the reward function by mutating a single component to enhance the agent’s driving capability further. Some helpful tips for analyzing the reward components: If the values for a certain reward component are nearly identical throughout, then this means the training is not able to optimize this component as it is written. You may consider: 1. Rescaling it to a proper range or the value of its temperature parameter 2. Re-writing the reward component 3. Discarding the reward component
@@
The mutation process involves the following steps: 1. First, select a reward component based on the given guidelines. 2. Next, clearly explain how you intend to mutate the selected component and your rationale behind improving the driving performance. This could involve adjusting its scale, rewriting its formula, or other modifications. 3. Finally, write the mutated reward function code.
@@
The output of the reward function should consist of two items: 1. The total reward, 2. A dictionary of each individual reward component.